Show that the notation {Xi} i€I implicitly involves the notion of function.


1
Expert's answer
2021-06-25T05:37:06-0400

If the equation F(x,y) =0 is written out in full terms of the coordinates, it will look like below

"F_1(x_1, ...,x_i, y_1, ...,y_i)=0"

"F_2(x_1, ...,x_i, y_1, ...,y_i)=0"

"."

"."

"."

"F_i(x_1, ...,x_i, y_1, ...,y_i)=0."

If "R^{i+1}\to R^i" then dxF is the differential of the function Ri+1 to Ri obtained by fixing and

∂F/∂x is the corresponding Jacobi matrix , thus

"\frac{∂F}{∂ x} = \begin{bmatrix}
 \frac{∂F_1}{∂ x_1} & \frac{∂F_1}{∂ x_i} \\
 \frac{∂Fi}{∂ x_1} & \frac{∂F_i}{∂ x_i} 
\end{bmatrix}"

Letting "R^{i+1}\to R^i" be Ci at a point (a,b) with F(a,b)=0. If dxF(a,b) is invertible , there are positive numbers € and 𝛿

a) If "|x-a|< \delta" then there is one and only one point "y= \varphi(x)" satisfying "|x-b|< \u20ac" and F(x,y)=0